# New polarimetric study of the galactic X-ray burster GX 13$+$1

### v1

@@
Weakly magnetized neutron stars (WMNS) are complicated sources with challenging phenomenology. For decades, they have been studied via spectrometry and timing. It has been established that the spectrum of WMNSs consists of several components traditionally associated with the accretion disk, the boundary or spreading layer, and the wind and their interactions with each other. Since 2022, WMNSs have been actively observed with the Imaging X-ray Polarimetry Explorer (IXPE). Polarimetric studies provided new information about the behavior and geometry of these sources. One of the most enigmatic sources of the class, galactic X-ray burster GX 13$+$1 was first observed with IXPE in October 2023. A strongly variable polarization at the level 2–5% was detected with the source showing a rotation of the polarization angle (PA) that hinted towards the misalignment within the system. The second observation was performed in February 2024 with a complementary observation by Swift/XRT. IXPE measured an overall polarization degree (PD) of 2.5% and the PA of 24°, and the Swift/XRT data helped us evaluate the galactic absorption and fit the continuum. Here we study the similarities and differences between the polarimetric properties of the source during the two observations. We confirm the expectation of the misalignment in the system and the assignment of the harder component to the boundary layer. We emphasize the importance of the wind in the system. We note the difference in the variation of polarimetric properties with energy and with time.
@@ 1 Introduction
Although weakly magnetized neutron stars (WMNS) do not show pulsation or strong orbital variation, they present a complex phenomenology that requires further studies and observations to be explained. The broad class of WMNSs includes sources with a magnetic field of the order of $10^{7}-10^{9}$ G. These neutron stars accrete matter from the companion via the Roche lobe, and if the accretion rate is high enough to overpower the relatively weak magnetic field, the matter from the accretion disk falls directly onto the equator of the neutron star.
@@
Non-pulsating weakly magnetized neutron stars are traditionally classified into two main groups: Z and atolls, on the basis of the patterns they show on the color-color diagram (CCD) or hardness-intensity diagram (HID). Although the classification is completely phenomenological, objects of these two classes have many other characteristic behavioral properties. For instance, atolls are known to be less bright than the Z-sources: typical luminosity of the atolls is in the range of $10^{36}-10^{37}$ erg s-1, while Z-sources are among the brightest objects in the X-ray sky with luminosities on the order of $10^{38}$ erg s-1. However, many features are known to be present in both classes: for example, quasi-periodic oscillations (QPOs) in the Hz and kHz ranges were observed from both Z and atolls sources (van der Klis, 1989, 2000). Sources of both classes are observed in both hard and soft states, and the main components of their spectrum are expected to be the same. The spectrum is generally well described by the two main components (see e.g. Revnivtsev et al., 2013). The soft component can be approximated by the multicolor black body emission of temperature $<1$ keV associated with the accretion disk. The harder component most probably comes from Comptonization in a relatively cool plasma (2–3 keV), either a boundary layer between the accretion disk and the surface of the neutron star (Shakura & Sunyaev, 1988), or the spreading layer at the neutron star surface (Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006). Additional features known to the spectrum are the broad Fe emission line associated with the reflection of the spreading layer emission from the accretion disk and sharp absorption lines coming from the absorption of the emission in the wind above the disk.
@@
However, the geometry of the WMNS cannot be fully understood from spectroscopic and time-resolved observations only. Information obtained from variability, and specifically from the QPOs, revealed that the strongest variability is associated with a harder spectral component (Gilfanov et al., 2003; Revnivtsev & Gilfanov, 2006; Revnivtsev et al., 2013) originating close to the neutron star surface. Polarimetry brought a new perspective to the problem. For instance, it became possible to distinguish between the boundary and spreading layer emission based on the polarimetric properties of the emission, while spectroscopically they are identical. The boundary layer is expected to have a polarization similar to that of the accretion disk, with the polarization vector lying in the plane of the disk (Chandrasekhar, 1960; Loktev et al., 2022), while the polarization of the spreading layer is expected to be perpendicular to the accretion disk (Farinelli et al., 2024). For example, the polarization angle (PA) of the bright Z-source Cyg X-2 (Farinelli et al., 2023) turned out to be consistent with the position angle of the radio jet (which is likely perpendicular to the accretion disk), implying the origin of the polarization in the spreading layer or from scattering in the accretion disk wind (Tomaru et al. 2024; Nitindala et al., in prep.). In La Monaca et al. (2024), on the other hand, a significant difference between the jet position angle and the PA for Sco X-1 has been observed which implies a more complicated geometry of the source. In the peculiar case of Cir X-1 (Rankin et al., 2024), the variability of the PA with time and hardness supported the idea of a misalignment between the angular momentum of the neutron star and the binary orbital axis. In Z-sources XTE J1701--462 (Cocchi et al., 2023) and GX 5--1 (Fabiani et al., 2023), a significant variation of the polarization between observations in the hard and soft states suggested the change in geometry between states. In atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820--303 (Di Marco et al., 2023a), and 4U 1624--49 (Saade et al., 2024), the strong dependency of the PD on energy supported the separation of the spectrum into two components with the different PAs. These discoveries supported the study of the WMNSs and established the capabilities of polarimetric observations of these sources.
@@
Galactic X-ray burster GX 13$+$1 was first observed by IXPE on 2023 October 17–19 (Bobrikova et al., 2024, hereafter B24). This source is known for the strong wind above the large disk being responsible for a number of absorption lines (Díaz Trigo et al., 2012), the brightness of $0.5L_{\rm Edd}$ (D’Aí et al., 2014), and an orbital period of 24.5 days, which is very long for a low-mass X-ray binary (Corbet et al., 2010). GX 13$+$1 presents a classification challenge: it is still debatable whether the source is a Z (see e.g. Saavedra et al., 2023) or an atoll (see e.g. Schnerr et al., 2003).
@@
The remainder of the paper is structured as follows. We introduce the observations performed with IXPE and Swift/XRT observatories in Sect. 2. In Sect. 3 we present the data analysis. We provide possible interpretations of the results in Sect. 4 and a summary in Sect. 5.
@@ 2.1 IXPE
IXPE is a joint mission of NASA and the Italian Space Agency (ASI) that was launched on 9 December 2021. The observatory comprises three grazing incidence mirror assembly modules, each having a detector unit (DU), hosting an X-ray detector sensitive to polarization, the Gas Pixel Detector (GPD), in the focal plane (Costa et al., 2001; Baldini et al., 2021). A description of the mission and of the instrument on board is given in Soffitta et al. (2021), and Weisskopf et al. (2022).
IXPE observed GX 13$+$1 two times, the first one performed in October 2023, reported in B24, and the present standing since 25 February 2024 at 13:34 UTC until 27 February at 12:59 UTC for a total exposure time of about 90 ks in each DU (see Table 1 and the light curve in Fig. 1). Data have been processed for model-independent polarimetric analysis using the ixpeobssim package version 30.2.1 (Baldini et al., 2022). Spectral and spectro-polarimetric analysis was performed using HEASoft version 6.33 and the standard FTOOLS (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc) with the IXPE CALDB released on 28 February 2024. Source photons were selected in a circular region with a radius of 100″  centered at the source position. As recommended by Di Marco et al. (2023b), given the brightness of the source, the background was not subtracted. The unweighted analysis is performed with ixpeobssim, while the weighted analysis is adopted for the spectropolarimetric analysis as suggested by Di Marco et al. (2022a).
Table 1: List of IXPE and Swift/XRT observations: Observation ID, start, stop, and the exposure times are reported.
Observatory ObsID Start – Stop Exposure Time (s)
Swift/XRT-XRT 00036688042 2024-02-26T05:05–2024-02-26T05:27 1366
@@
IXPE 03001101 2024-02-25T13:34–2024-02-27T12:59 90670
90819
90799
Figure 1: The IXPE light curve obtained combining the 3 DUs (top panel) and the corresponding hardness ratio as a function of time (bottom panel). The data are binned in 200 s. The shaded regions mark the Swift/XRT observation periods. Dashed vertical lines illustrate the separation of the observation into five equal time bins of 9.5 h.
@@ 2.2 Swift/XRT
The Neil Gehrels Swift Observatory (Gehrels et al., 2004) carrying three instruments on board, enables the most detailed observations of gamma-ray bursts to date; the X-ray Telescope (XRT) is one of the instruments on board based on a sensitive, flexible, autonomous X-ray CCD imaging spectrometer. Swift/XRT coordinated observations with IXPE for GX 13$+$1 were performed; they are used to monitor the status of the source and obtain spectral information during IXPE observation. Given the source brightness, the Swift/XRT observations were performed in Windowed Timing mode (WT). Ten Swift/XRT pointings covered the IXPE observation, as reported in Table 1 and Fig. 1. Swift/XRT data were extracted using HEASoft v6.33 and standard FTOOLS Nasa High Energy Astrophysics Science Archive Research Center (2014) (Heasarc). Source and background extractions were performed using Swift/XRT imaging capabilities in an annulus with an inner radius of about 5″ and an outer radius of about 60″ to avoid possible pile-up effects (Romano et al., 2006). The Swift/XRT data have been fitted in the energy band 0.7–8 keV to have sufficient statistics in the spectral fits and grouped to have at least 50 counts bin-1. In the analysis, the response matrices released in the HEASARC CALDB on 2023 July 25 have been applied.
@@ 3 Data analysis
IXPE observed GX 13$+$1 for $\sim$90 ks. During the observation, as shown in Fig. 1 (top panel), the source rate was rather stable, with a slight increase at the end of the observation. The source hardness in the IXPE data appears to have no strong variations, remaining almost constant throughout the observation; this is in contrast to the previous IXPE observation, reported in B24, where a dip corresponding to a slight hardening was present. In Fig. 2, the hardness intensity diagram obtained from the IXPE data is reported and shows that hardness did not change by more than 10%.
Swift/XRT observations, covering with short snapshots the IXPE observation, allow us to monitor the source spectral state. GX 13$+$1 studies in the past showed a spectrum that can be well described, as typical in the atoll and Z sources, by a soft diskbb component describing disk emission and Comptonization emission from the spreading/boundary layer that can be described by a bbodyrad component (Díaz Trigo et al., 2012). Better modeling of the hard Comptonization component, e.g., applying compTT or compTB, is not possible due to the limited IXPE and Swift/XRT energy bands. In the literature, XMM–Newton spectra of GX 13$+$1 are reported showing wind and reflection features (Díaz Trigo et al., 2012), but neither IXPE nor Swift/XRT allow us to study these features, so in the following the simple spectral model tbabs*(diskbb+bbodyrad) will be applied.
@@ 3.1 Spectroscopy
Given the better spectral capabilities, we used Swift/XRT data as a driver for spectral modeling and to monitor spectral variations. The analysis is carried out in the energy band 0.7–8 keV, best-fit values are reported for each observation in Table 2 with uncertainties at the 68% confidence level (CL), and ObservationIDs are ordered by the observation date to show the time evolution of the spectral model.
Table 2: Best-fit parameters for the spectral model TBabs*(diskbb+bbodyrad) as obtained by the different Swift/XRT spectra.
@@
$N_{\text{H}}$ (1022 cm2) $kT_{\text{in}}$ (keV) norm $kT_{\text{bb}}$ (keV) norm (10-9 erg cm-2 s-1)
@@
 [^1] The fits are performed in the energy band 0.7–8.0 keV. Errors correspond to the 68% CL. The reported flux has a typical uncertainty of the order of $0.3\times 10^{-9}$. The hardness is defined as in Eq. (1).
In the spectral analysis, the Swift/XRT data are fitted with the model: tbabs*(diskbb+bbodyrad). In the different pointings, it is possible to observe that the absorption is in average $\sim 4.7\times 10^{22}$ cm-2, while the disk temperature ranges between 0.7–1.1 keV, about the bbodyrad component – which in our case is used to describe a Comptonization emission from the boundary layer – the temperature ranges in the $\sim 1.3-1.8$ keV. The flux is in the range $\sim$(4–6)×$\times 10^{-9}$ erg cm-2 s-1 except in the last pointing where it reaches the highest value of $\sim 8\times 10^{-9}$ erg cm-2 s-1. The hardness, reported in Table 2 is defined as
$\text{Hardness}=\frac{\text{Flux}_{0.7-3\leavevmode\nobreak\ \text{keV}}-\text{Flux}_{3-8\leavevmode\nobreak\ \text{keV}}}{\text{Flux}_{0.7-8\leavevmode\nobreak\ \text{keV}}}$ (1)
and the spectral models report an almost constant hardness value except for the case of ObservationID 00036688045 where the flux is at a minimum and there is a slight hardening, but this Swift/XRT observation (third snapshot) is performed mainly during an IXPE occultation period. Thus, the IXPE data are not affected by this slight hardening, while the other snapshots confirm the IXPE result of an almost constant hardness ratio during observation. In the following analysis, for the spectro-polarimetric study, this ObservationID will be excluded, while the other ones will be used in the joint fit with the IXPE data.
Taking into account the whole IXPE observation, and the Swift/XRT spectra, except the ObservationID 00036688045, we obtain a good joint spectrum model. The best-fit parameters are reported in Table 3 and Fig. 3. To take into account the calibration uncertainties in the IXPE data, (see, e.g., Di Marco et al., 2022b), we left the gain slope and offset of the IXPE-DUs free to vary.
Table 3: Best-fit parameters of the model const*tbabs*(diskbb+bbodyrad) for the joint Swift/XRT and IXPE spectra.
@@
Figure 3: Joint spectral fit Swift/XRT and IXPE in the 1–8 keV energy range. In the top panel the spectra of each telescope are reported in $EF_{E}$ units, in the bottom panel the residuals are shown.
The joint fit confirms the result expected for the average model from several Swift/XRT snapshots. In the following for the spectro-polarimetric analysis, the spectral model will be frozen to the present one. The IXPE gain correction is at level of 0.95 for the slope in all the three DUs and the offset in the range 70–100 eV.

### v2

@@
Weakly magnetized neutron stars (WMNSs) are complex astrophysical objects with challenging phenomenology. For decades, they have been studied via spectrometry and timing analyses. It is well established that the spectrum of WMNSs consists of several components traditionally associated with the accretion disk, the boundary or spreading layer, and the wind, along with their interactions. Since 2022, WMNSs have been actively observed using the Imaging X-ray Polarimetry Explorer (IXPE). Polarimetric studies have provided new information about the behavior and geometry of these sources. One of the most enigmatic sources in this category, the galactic X-ray burster GX 13$+$1, was first observed with IXPE in October 2023. A highly variable polarization at levels of 2–5% was detected, with the source showing a rotation of the polarization angle (PA), suggestive of misalignment within the system. A second observation was performed in February 2024, complemented by observations from Swift/XRT. IXPE measured an overall polarization degree (PD) of 2.5% and a PA of 24°, while Swift/XRT data helped us evaluate the galactic absorption and fit the continuum. Here, we study the similarities and differences in the polarimetric properties of the source during the two observations. Our findings confirm the expected misalignment in the system and the assignment of the harder component to the boundary layer. We also emphasize the significance of the wind in the system. Additionally, we observe notable differences in the variation of polarimetric properties with energy and over time.
@@ 1 Introduction
Although weakly magnetized neutron stars (WMNSs) do not show pulsation or strong orbital variation, they present a complex phenomenology that requires further studies and observations to be explained. The broad class of WMNSs includes sources with a magnetic field on the order of $10^{7}-10^{9}$ G. These neutron stars accrete matter from their companions via the Roche lobe. If the accretion rate is high enough to overpower the relatively weak magnetic field, the matter from the accretion disk falls directly onto the equator of the neutron star.
@@
Non-pulsating WMNSs are traditionally classified into two main groups, Z and atolls, based on the patterns they show on the color-color diagram (CCD) or hardness-intensity diagram (HID). Although the classification is completely phenomenological, objects of these two classes have many other characteristic behavioral properties. For instance, atolls are known to be less bright than the Z sources: the typical luminosity of the atolls is in the range of $10^{36}-10^{37}$ erg s-1, while Z sources are among the brightest objects in the X-ray sky, with luminosities on the order of $10^{38}$ erg s-1. However, many features are known to be present in both classes. For example, quasi-periodic oscillations (QPOs) in the hertz and kilohertz ranges were observed from both Z and atoll sources (van der Klis, 1989, 2000). Sources of both classes are observed in both hard and soft states, and the main components of their spectrum are expected to be the same. The spectrum is generally well described by the two main components (see, e.g., Revnivtsev et al., 2013). The soft component can be approximated by the multicolor blackbody emission of temperature $<1$ keV associated with the accretion disk. The harder component is probably caused by Comptonization in a relatively cool plasma (2–3 keV), either a boundary layer between the accretion disk and the surface of the neutron star (Shakura & Sunyaev, 1988) or the spreading layer at the neutron star surface (Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006). Additional features known to the spectrum are the broad Fe emission line associated with the reflection of the spreading layer emission from the accretion disk and sharp absorption lines caused by the absorption of the emission in the wind above the disk.
@@
However, the geometry of the WMNSs cannot be fully understood from spectroscopic and time-resolved observations alone. Information obtained from variability, and specifically from the QPOs, revealed that the strongest variability is associated with a harder spectral component (Gilfanov et al., 2003; Revnivtsev & Gilfanov, 2006; Revnivtsev et al., 2013) originating close to the neutron star surface. Polarimetry brought a new perspective to the problem. For instance, it became possible to distinguish between the boundary and spreading layer emission based on the polarimetric properties of the emission, while spectroscopically, they are identical. The boundary layer is expected to have a polarization similar to that of the accretion disk, with the polarization vector lying in the plane of the disk (Chandrasekhar, 1960; Loktev et al., 2022), while the polarization of the spreading layer is expected to be perpendicular to the accretion disk (Farinelli et al., 2024). For example, the polarization angle (PA) of the bright Z-source Cyg X-2 (Farinelli et al., 2023) turned out to be consistent with the position angle of the radio jet (which is likely perpendicular to the accretion disk), implying the polarization originated in the spreading layer or from scattering in the accretion disk wind (Tomaru et al. 2024; Nitindala et al., in prep.). In La Monaca et al. (2024), on the other hand, a significant difference between the jet position angle and the PA for Sco X-1 has been observed, which implies a more complicated geometry of the source. In the peculiar case of Cir X-1 (Rankin et al., 2024), the variability of the PA with time and hardness supports the idea of a misalignment between the angular momentum of the neutron star and the binary orbital axis. In Z-sources XTE J1701--462 (Cocchi et al., 2023) and GX 5--1 (Fabiani et al., 2024), a significant variation in the polarization between observations in the hard and soft states suggested a change in geometry between states. In atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820--303 (Di Marco et al., 2023), and 4U 1624--49 (Saade et al., 2024), the strong dependence of the polarization degree (PD) on energy supports the separation of the spectrum into two components with the different PAs. These discoveries demonstrate the usefulness of using polarimetric observations to study these sources.
@@
The galactic X-ray burster GX 13$+$1 was first observed by Imaging X-ray Polarimetry Explorer (IXPE) on 2023 October 17–19 (Bobrikova et al., 2024, hereafter B24). This source is known for the strong wind above the large disk being responsible for several absorption lines (Díaz Trigo et al., 2012), a brightness of $0.5L_{\rm Edd}$ (D’Aí et al., 2014), and an orbital period of 24.5 d, which is very long for a low-mass X-ray binary (Corbet et al., 2010). GX 13$+$1 presents a classification challenge: it is still debatable whether it is a Z (see, e.g., Saavedra et al., 2023) or an atoll source (see, e.g., Schnerr et al., 2003).
@@
The remainder of the paper is structured as follows: we introduce the observations performed with the IXPE and Swift/XRT observatories in Sect. 2. In Sect. 3, we present our data analysis. We provide possible interpretations of the results in Sect. 4 and a summary in Sect. 5.
@@ 2.1 IXPE
IXPE is a joint mission of the National Aeronautics and Space Administration (NASA) and the Italian Space Agency (Agenzia Spaziale Italiana, ASI) that was launched on 2021 December 9. The observatory comprises three grazing-incidence mirror assembly modules, each of which has a polarization-sensitive X-ray detector unit (DU) in the focal plane (Costa et al., 2001; Baldini et al., 2021). A description of the mission and of the instrument on board is given in Soffitta et al. (2021) and Weisskopf et al. (2022).
IXPE observed GX 13$+$1 twice: the first observation was conducted in October 2023, as reported in B24, and the second took place from 2024 February 25, at 13:34 UTC to 2024 February 27, at 12:59 UTC, with a total exposure time of approximately 90 ks for each DU (see Table 1 and the light curve in Fig. 1). Data were processed for model-independent polarimetric analysis using the ixpeobssim package version 30.2.1 (Baldini et al., 2022). Spectral and spectropolarimetric analysis was performed using HEASoft version 6.33 and the standard FTOOLS (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc), with the IXPE Calibration Database (CALDB) that was released on 2024 February 28. Source photons were selected in a circular region with a radius of 100″ centered at the source position. As recommended by DiMarco_2023, given the brightness of the source, the background was not subtracted. The unweighted analysis was performed with ixpeobssim, while the weighted analysis was adopted for the spectropolarimetric analysis, as suggested by Di Marco et al. (2022a).
Table 1: List of IXPE and Swift/XRT observations.
Observatory ObsID Start – Stop Instrument Exposure Time (s)
Swift 00036688042 2024-02-26T05:05–2024-02-26T05:27 XRT 1366
@@
IXPE 03001101 2024-02-25T13:34–2024-02-27T12:59 DU1 90670
DU2 90819
DU3 90799
Figure 1: IXPE light curve obtained by combining the three DUs (top panel) and the corresponding hardness ratio as a function of time (bottom panel). The data are binned in 200 s. Shaded regions indicate the Swift/XRT observation periods. Dashed vertical lines indicate the separation of the observation into five equal 9.5 h time bins.
@@ 2.2 Swift/XRT
Carrying three instruments on board, the Neil Gehrels Swift Observatory (Gehrels et al., 2004) enables the most detailed observations of gamma-ray bursts to date; the X-ray Telescope (XRT), one of the instruments on board, is based on a sensitive, flexible, autonomous X-ray CCD imaging spectrometer. Swift/XRT-coordinated observations with IXPE for GX 13$+$1 were performed, and used to monitor the status of the source and obtain spectral information during IXPE observations. Given the source brightness, the Swift/XRT observations were performed in windowed timing (WT) mode. Ten Swift/XRT pointings covered the IXPE observation, as reported in Table 1 and Fig. 1. Swift/XRT data were extracted using HEASoft v6.33 and standard ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc). Source and background extractions were performed using Swift/XRT imaging capabilities in an annulus with an inner radius of about 5″ and an outer radius of about 60″ to avoid possible pile-up effects (Romano et al., 2006). The Swift/XRT data were fitted in the energy band 0.7–8 keV to ensure sufficient statistics in the spectral fits and grouped to have at least 50 counts bin-1. Response matrices released in the HEASARC CALDB on 2023 July 25, were applied in this analysis.
@@ 3 Data analysis
IXPE observed GX 13$+$1 for $\sim$90 ks. During the observation, as shown in Fig. 1 (top panel), the source rate was rather stable, with a slight increase at the end of the observation. The source hardness in the IXPE data appears to have no strong variations, showing only a slight decrease with the count rate increase toward the end of the observation; this contrasts with the previous IXPE observation reported in B24, where a dip corresponding to a slight hardening was present. Figure 2, which reports the HID obtained from the IXPE data, shows that hardness did not change by more than 10%.
Consisting of short snapshots during the IXPE observation, the Swift/XRT pointings enable us to monitor the spectral state of the source. Previous studies of GX 13$+$1 have shown that its spectrum can be well described, as is typical for atoll and Z sources, by a soft diskbb component representing disk emission and a bbodyrad component representing Comptonization emission from the spreading and/or boundary layer (Díaz Trigo et al., 2012). More detailed modeling of the hard Comptonization component, such as using compTT or compTB, is not possible due to the limited energy bands of IXPE and Swift/XRT. Literature on XMM–Newton spectra of GX 13$+$1 reports wind and reflection features (Díaz Trigo et al., 2012), but neither IXPE nor Swift/XRT allows us to study these features. Therefore, in the following analysis, we will apply the simplified spectral model, tbabs*(diskbb+bbodyrad).
@@ 3.1 Spectroscopy
Given its superior spectral capabilities, we used Swift/XRT data as the primary source for spectral modeling and to monitor spectral variations. The analysis was conducted in the 0.7–8 keV energy band, with best-fit values reported for each observation in Table 2, including uncertainties at the 68% confidence level (CL). Observation IDs are ordered by observation date to illustrate the time evolution of the spectral model.
Table 2: Best-fit parameters for the TBabs*(diskbb+bbodyrad) spectral model as obtained by the different Swift/XRT spectra.
@@
$N_{\text{H}}$ (1022 cm-2) $kT_{\text{in}}$ (keV) norm $kT_{\text{bb}}$ (keV) norm (10-9 erg cm-2 s-1)
@@
 [^1] The fits are performed in the energy band 0.7–8.0 keV. Errors correspond to the 68% CL. The reported flux has a typical uncertainty on the order of $0.3\times 10^{-9}$. The hardness is defined as in Eq. (1).
In the spectral analysis, the Swift/XRT data were fitted using the tbabs*(diskbb+bbodyrad)model. Across the different pointings, the absorption column density was on average $\sim 4.7\times 10^{22}$ cm-2, while the disk temperature ranged between 0.7 and 1.1 keV. Regarding the bbodyrad component – which in our case is used to describe a Comptonization emission from the boundary layer – the temperature ranged from $\sim 1.3-1.8$ keV. The flux was typically in the range of $\sim$(4–6)×$\times 10^{-9}$ erg cm-2 s-1, except in the final pointing, where it reached the highest value of $\sim 8\times 10^{-9}$ erg cm-2 s-1. The hardness (see Table 2) defined as
$\text{Hardness}=\frac{\text{Flux}_{3-8\leavevmode\nobreak\ \text{keV}}-\text{Flux}_{0.7-3\leavevmode\nobreak\ \text{keV}}}{\text{Flux}_{0.7-8\leavevmode\nobreak\ \text{keV}}}$ (1)
has an almost constant value, except for Observation ID 00036688045, where the flux was minimal and the hardness maximal. However, this Swift/XRT observation (third snapshot) was conducted mainly during an IXPE occultation period. Thus, the IXPE data are not affected by this slight hardening, and the other snapshots confirm the IXPE result of an almost constant hardness ratio during the observation. In the following analysis for the spectropolarimetric study, this Observation ID is excluded, and the others are used in the joint fit with the IXPE data.
Considering the entire IXPE observation and the Swift/XRT spectra, except the Observation ID 00036688045, we obtain a good joint spectrum model. The best-fit parameters are reported in Table 3 and Fig. 3. To account for the calibration uncertainties in the IXPE data (see, e.g., Di Marco et al., 2022b), we allowed the gain slope and offset of the IXPE DUs to vary freely.
Table 3: Best-fit parameters of the const*tbabs*(diskbb+bbodyrad) model for the joint Swift/XRT and IXPE spectra.
@@
Figure 3: Joint spectral fit of Swift/XRT and IXPE data in the 1–8 keV energy range. In the top panel, the spectra of each telescope are reported in $EF_{E}$ units; residuals are shown in the bottom panel.
The joint fit confirms the expected result for the average model from several Swift/XRT snapshots. In the following spectropolarimetric analysis, the spectral model is fixed to the present one. The IXPE gain correction is at a level of 0.95 for the slope in all three DUs, with the offset in the range of 70–100 eV.